OpenStudy (anonymous):

HELP ME PLEASE!!! Find the roots of the polynomial equation. 2x^3 + 2x^2 – 19x + 20 = 0 (1 point)

OpenStudy (anonymous):

A. \[\frac{ 3+i }{ 2}, \frac{ 3-i }{ 2 }, -4\] B. \[\frac{ -3+i }{ 2 }, \frac{ -3-i }{ 2}, 4\] C. \[\frac{ -3+i }{ 2}, \frac{ -3-i }{ 2 },-4\] D. \[\frac{ 3+2i }{ 2}, \frac{ 3-2i }{ 2}, 4\]

OpenStudy (anonymous):

@amistre64 @agent0smith @Ashleyisakitty @grant330sims @OtonoGold @austinL anyone help ?

OpenStudy (otonogold):

f(x) = 2x^3+2x^2-19x+20 Using rational root theorem to test root, you can find f(-4) = 0. So, one root is x = -4. Now you can reduce the degree by synthetic division, -4 | 2 2 -19 20 .......-8 24 -20 _________________ ....2 -6..5 | 0 Solve 2x^2 - 6x + 5 = 0 x = (1/2)[3 +/- i] <==two imaginary roots

OpenStudy (otonogold):

The answer is A. Im a 100 % sure

OpenStudy (anonymous):

thank you :)

OpenStudy (otonogold):

No problem :)